# Retrieval-Enhanced Machine Learning: Synthesis and Opportunities

### v1

@@ 5.2. Consumption
In REML, the predictive model is presented with one or more documents. The effectiveness of $f_{\theta}$ is influenced by the consumption of the presented documents. Ideally, $f_{\theta}$ would consume all the documents simultaneously, yet our systems are computationally limited; hence, the granularity of consumption is typically limited to a subset of the presented documents. Depending on the task, different consumption algorithms may prove varying in utility—some algorithms are used for extraction and others for on-the-fly updates of the predictive model parameters. In contrast, decoding algorithms, such as beam search (Freitag and Al-Onaizan, 2017) or nucleus sampling (Holtzman et al., 2020), provide ways to decode effective outputs given the presented documents and can incorporate verification for improved effectiveness. There are additional concerns during consumption, such as efficiency (De Jong et al., 2023; Hofstätter et al., 2023) and attribution (Gao et al., 2023b; Schuster et al., 2024; Asai et al., 2024; Menick et al., 2022), that provide further utility.
@@ 5.2.1. Consumption at different granularities of retrieval items
  • Joint: When the prediction has a sufficiently flexible context representation, then multiple retrieval items can be passed simultaneously in a single inference procedure (Izacard and Grave, 2021b; Lewis et al., 2020a). This is potentially richer than the ensemble approach since each retrieved item is aware of the others. Due to computational limits, only a few retrieval items are able to be processed this way and the ensemble-approach is relatively more scalable.
@@
These paradigms are atomic functions that can be composed to form more complex operations. For instance, given lists of retrieved items, denoted as $r_{0},r_{1},r_{2},r_{3}$, the atomic functions with different granularities can be composed as the following equation:
@@ 5.2.2. Consumption algorithms
  • Extractive: The predictive model is limited to extracting exact information from the retrieved item, e.g., a span of text from a retrieve passage to answer a question (Khandelwal et al., 2020; Lan et al., 2023). This can be achieved through pointer networks (Vinyals et al., 2015), constrained decoding (Hokamp and Liu, 2017; Hu et al., 2019; Post and Vilar, 2018), and other similar techniques.
@@ 5.2.3. Decoding algorithms
Independent of the consumption algorithm, there are different decoding algorithms that can be done to predict a high quality output in REML. Broadly, decoding algorithms explored thus far fall into the following categories:
@@
  • Retrieval-enhanced: Search algorithms like beam search will consider both the model output and the retrieved items when scoring candidate predictions (Lewis et al., 2020a; Asai et al., 2024). This should penalize spurious associations between query and retrieved item.
@@ 5.2.5. Attribution and other extensions
Advanced applications of REML will augment the predictive model output space to incorporate REML-specific information. The most common instance of this is probably to support attribution, so that each part of the output can be traced back to the relevant retrieved item (Gao et al., 2023b; Schuster et al., 2024; Asai et al., 2024; Menick et al., 2022). Other cases involve in-line verification or calls to external tools that would not be easily possible without incorporating the appropriate retrieved item (Asai et al., 2024).
Consumption Granularities (§5.2.1) $k$NN-LM (Khandelwal et al., 2020) Single + Ensemble Probabilities are computed for each retrieved item independently then combined. RePLUG (Shi et al., 2024) Single + Ensemble Probabilities are computed for each retrieved item independently then combined. RAG (Lewis et al., 2020a) Joint Retrieved items are concatenated before consumption. FiD (Izacard and Grave, 2021b) Joint Retrieved items are concatenated in the decoder during consumption. FLARE (Jiang et al., 2023) Multi-round Potentially retrieves new items as generation progresses. Consumption Algorithms (§5.2.2) kNN-LM (Khandelwal et al., 2020) Extractive A single word is selected from the retrieved context. CoG (Lan et al., 2023) Extractive An extension of kNN-LM that can output both words and phrases. CBR (Das et al., 2020) Analogical Incorporate knowledge graphs into neural models in spirt of case based reasoning. Dynamic L2M (Drozdov et al., 2023) Analogical Retrieves demonstrations for few-shot prompting. RePLUG (Shi et al., 2024) Contextual Uses retrieved items in the context. SPALM (Yogatama et al., 2021) Latent Incorporates retrieved items into the hidden state. Decoding Algorithms (§5.2.3) kNN-LM (Khandelwal et al., 2020) Text-only Retrieval probability is ignored for next word prediction. RAG (Lewis et al., 2020a) Retrieval-enhanced Retrieval probability is incorporated in beam search. Self-RAG (Asai et al., 2024) Retrieval-enhanced Critic probability is incorporated in beam search. FLARE (Jiang et al., 2023) Verification-based Discards low confidence continuations, triggering retrieval. Consumption Efficiency (§5.2.4) LUMEN (De Jong et al., 2023) Precompute Partially computes passage representations offline. Attribution and Extensions (§5.2.5) ALCE (Gao et al., 2023b) Attribution End of sentence. Multi-source. SemQA (Schuster et al., 2024) Attribution Mid-sentence. Multi-source. Self-RAG (Asai et al., 2024) Attribution End of sentence. Single source. Verification Outputs a special token indicating whether to use document.
@@ 6. Storing
Storing refers to how retrievable items are represented and indexed. how do we define $\mathcal{D}$? how does $g_{\omega}$ construct each $d\in\mathcal{D}$?
Storing is one of the optional yet crucial properties of REML models and refers to how retrievable items are saved, represented, and indexed. The storage components can be categorized into coupled and decoupled storage. If at least one external memory is optimized jointly with the predictive model, we call the architecture has coupled storage. In the coupled storage architecture, contents can be populated to the storage online, updated alongside a predictive model, e.g., Neural Turing Machines (NTM) (Graves et al., 2014, 2016) and REALM (Guu et al., 2020). On the other hand, if all external storage are from off-the-shelf system, and the contents are populated offline, we call the architecture has decoupled storage, e.g., kNN-LM (Khandelwal et al., 2020) and ED2LM (Hui et al., 2022b).
In its simplest form, REML systems will operate with decoupled storage where the retrieval model is implemented independent of the predictive model. Since entries in the storage are populated offline and many off-the-shelf retrieval models are readily available, it is relatively convenient to construct the decoupled storage. On the other hand, many advanced REML systems operate with coupled storage where the retrieval model is directly influenced by the predictive model. Here, the entries in the storage are populated online and updated alongside the predictive models. In this section, we will describe storage operations and challenges associated with the coupled and decoupled storage.
@@ 6.1. Primary Storage Operations
In general, there are three types of operations that each storage system must support to be effectively utilized in REML systems: Address Generation, Read, and write. The storage operations will typically be used in three scenarios: 1) $f_{\theta}$ needs to incorporate historical context in its prediction, e.g., long context language models; 2) $f_{\theta}$ conducts various types of online learning using recent/past interactions, e.g., experience replay in reinforcement learning and language agents; and 3) $f_{\theta}$ is a memory network-like architecture (Weston et al., 2015) where retrieval is an abstract process and everything written to or read from the storage is controlled by the network which is the service of the objective being optimized.
@@ 6.1.3. Write
where payload can be a form of vector or raw representation, which can be preprocessed by some functions, similar to the functions defined in Section 5, before being stored in the storage (Hui et al., 2022b).
@@ 6.2.1. Storage Construction
In most cases, a REML system will initialize its storage by processing an entire retrieval corpus. This can be done offline before training, after training before inference, and throughout training as needed. Storage construction is well documented and studied, and the initial storage construction is essentially a series of write and address generation operations. Storage can be constructed as a key-value structure where retrieval space $\mathcal{D}$ can be defined as:
@@
where $transform_{k}$ is a key representation function that can take each entry in the corpus or an input instance $x$. Similarly, $transform_{v}$ is a value representation function that can take each entry in the corpus or an input instance $x$ to generate a value in the storage. It is important to note that the $transform_{k}$ and $transform_{v}$ function can simply be the identity function, meaning it does not change the key and value at all. Table 4 describes how each paper constructed its storage offline and/or online. For example, in EMAT (Wu et al., 2022d), where the collection is pairs of questions and answers, the $transform_{k}$ function utilizes only the question in each pair to generate the key representation, and $transform_{v}$ function uses only the answer from each pair to generate the value representation.
@@ 6.2.2. Storage Management
Once the storage is initialized, both for optimal task completion and efficiency, there is a need to schedule the storage operations, and we dub this phase as storage management. Efficient usage of storage can be understood in terms of space and speed which come down to determining three decisions: when, what, and how to store.
@@
In the most scenarios, $g_{\omega}$ will pull information from the storage built at an initial phase based on the need of $f_{\theta}$. However, this setup can introduce storage staleness problem when $\omega$ changes during $g_{\omega}$ is being sequentially or jointly trained with $f_{\theta}$. Since $\mathcal{D}_{\text{key}}$ (retrieval space of keys in C) is constructed by $\omega$ during the training, it is necessary to refresh the storage as $\omega$ is updated; this was even mentioned from MemNN (Weston et al., 2015). This comes down to asking two questions: 1) how often to update, and 2) what portion of storage to update. For the first question, one can either synchronously (every training step) or asynchronously (every $T$ training step) update the storage (Asai et al., 2023a). For the second question, one can choose to update the entire storage, a subset of the storage, or refrain from updating the storage at all.
@@
This encompasses entry representation, e.g., index compression (Wu et al., 2022b; Rae et al., 2020; Martins et al., 2022) and quantization (Izacard et al., 2024) and architectural choice of the storage, e.g., key-value structure (Grave et al., 2017; Khandelwal et al., 2020; Wu et al., 2022c; Zhong et al., 2022; Min et al., 2023; Borgeaud et al., 2022; Yogatama et al., 2021; Alon et al., 2022; Hui et al., 2022b), where the compression and representation computations can happen incrementally by batch (Zamani et al., 2022).
@@ 6.3. Storage Types
Coupled Storage (§6.3.1) Key Value NTM (Graves et al., 2014, 2016) transformed output of $f_{\theta}$ (the same as the Key) MemNN (Weston et al., 2015) & MemN2N (Sukhbaatar et al., 2015) input feature embedding (the same as the Key) Dynamic Memory Network (Kumar et al., 2016) input feature embedding (the same as the Key) Neural Cache Model (Grave et al., 2017) hidden representation of RNN next word RUM (Chen et al., 2018) user-item embedding (the same as the Key) Transformer-XL (Dai et al., 2019) hidden representation of Transformer (the same as the Key) LongMem (Wang et al., 2023a) attention-key attention-value Memorizing Transformer (Wu et al., 2022c) attention-key attention-value MemTransformer (Burtsev et al., 2021) sequence of tokens tokens RPT (Rubin and Berant, 2023) token chunk embedding (the same as the Key) Unlimiformer (Bertsch et al., 2023) token chunk embedding (the same as the Key) MeMViT (Wu et al., 2022b) image frame embedding (the same as the Key) PFMN (Lee et al., 2018) image frame embedding (the same as the Key) REALM (Guu et al., 2020) document embedding (the same as the Key) REPLUG LSR (Shi et al., 2024) document embedding (the same as the Key) ATLAS (Izacard et al., 2024) document embedding (the same as the Key) EMAT (Wu et al., 2022d) question embedding answer embedding QAMAT (Chen et al., 2023c) question embedding question-answer embedding TRIME (Zhong et al., 2022) context (sequence of tokens) next token NPM (Min et al., 2023) token embedding token Reflexion (Shinn et al., 2023) NL self-reflection in NL (the same as the Key) CLIN (Majumder et al., 2024) self-reflection in NL (the same as the Key) ExpeL (Zhao et al., 2024) self-reflection in NL (the same as the Key) Generative Agent (Park et al., 2023) stream of experience in NL (the same as the Key) Voyager (Wang et al., 2023c) program description embedding program code MemPrompt (Madaan et al., 2022) NL question NL human feedback Decoupled Storage (§6.3.2) kNN-LM (Khandelwal et al., 2020) context embedding next token SPALM (Yogatama et al., 2021) context embedding next token RAG (Lewis et al., 2020b) document embedding (the same as the Key) RETOMATON (Alon et al., 2022) context embedding next token, pointer KIF (Fan et al., 2021) evidence embedding (multimodal) (the same as the Key) RETRO (Borgeaud et al., 2022) token chunk embedding token chunk ED2LM (Hui et al., 2022b) document embedding document REPLUG (Shi et al., 2024) document embedding (the same as the Key)
@@
In the literature, two types of storage architectures are identified in REML systems: Coupled Storage and Decoupled Storage. The following sections will elaborate on these architectures and introduce the characteristics associated with each type of storage.
@@ 6.3.1. Coupled Storage
Coupled storage is defined as a storage that can be updated online during training and inference of the predictive model and can be jointly optimized. Initial developments of the coupled storage enhancements were primarily led by Neural Turing Machine (NTM) (Graves et al., 2014, 2016) and Memory Network (MemNN) (Weston et al., 2015). They showed abstract operations, such as copying, recall, and sorting, to language reasoning tasks by leveraging external addressable storage, where contents in the storage of these architectures are dense vectors which can be readily used by $f_{\theta}$. We refer the readers to the original papers of the two models for deeper understanding of primary shape of REML models with coupled storage. There are a few notable characteristics about REML systems with coupled memory, including but not limited to:
@@
One of the primary concerns of the coupled storage arises when $g_{\omega}$ is being trained, making the storage stale. It is still an ongoing challenge in the community, but as mentioned in Section 6.2, there have been several techniques devised to circumvent this issue by answering how often to update (synchronous or asynchronous) and what portion of the storage to update (full or partial). Therefore, there are five different strategies including an avoidance of update. Synchronous full update is the simplistic approach to solve the staleness problem by updating the storage at every training step. It is attempted in a few research (Rubin and Berant, 2023; Bertsch et al., 2023), but its large computational overhead prevents it from being used in a practical setting (Izacard et al., 2024). Synchronous partial update can be done by selecting a batch of entries to update (Izacard et al., 2024). Depending on the applications, there can be various batch selection strategies, such as lexical similarity (Zhong et al., 2022) or in-document sampling (Min et al., 2023). Asynchronous full update is done by updating the full storage every $T$ training steps (Guu et al., 2020; Izacard et al., 2024; Shi et al., 2024; Wu et al., 2022d). This allows staleness in the index before it is updated again. For example, Wu et al. (2022d) freeze the storage at the beginning of each training epoch and only updates at the end of each epoch. As far as we know, there is little attempt on asynchronous partial update as it may degrade the training performance in larger margin. Alternatively, since index recreation is highly expensive, it is possible to ignore the staleness and avoid re-indexing with adequate strategies without compromising a performance in a large margin (Rae et al., 2016; Lewis et al., 2020b; Izacard et al., 2024; Guu et al., 2020; Wang et al., 2023a; Wu et al., 2022c).
@@
Another characteristic of the models equipped with coupled storage is that they can have a cold start problem, where the performance of $f_{\theta}$ is suboptimal before the storage is filled up with enough information. Most of the architectures that start with an empty storage such as language agents (LA) or long-context language models that process a long document in multiple training steps have this issue (Park et al., 2023; Majumder et al., 2024; Shinn et al., 2023; Zhao et al., 2024; Wang et al., 2023c; Wu et al., 2022c; Zhong et al., 2022). However, the cold start problem can be alleviated when the model can be adapted to a new task and the storage/experience built during the previous tasks is transferable to the new task (Majumder et al., 2024).
@@
Despite its disadvantages, coupled storage has seen significant development both theoretically and in various applications. It includes making REML systems with coupled storage end-to-end trainable (Sukhbaatar et al., 2015), capturing the position and temporality of language using episodic storage (Kumar et al., 2016), and scaling the storage through sparse access (Rae et al., 2016; Zaremba and Sutskever, 2016) and strategic storage management (Lample et al., 2019; Grave et al., 2017). These have led to applications in various domains, such as meta-learning (Santoro et al., 2016), sequential recommendation (Chen et al., 2018), video summarization and recognition (Lee et al., 2018; Wu et al., 2022b). In long context language modeling (Bertsch et al., 2023; Zhong et al., 2022), there have been approaches to solve the task through attention recurrence (Dai et al., 2019; Yogatama et al., 2021; Wu et al., 2022c; Wang et al., 2023a) and compression of hidden states (Rae et al., 2020; Martins et al., 2022; Wu et al., 2022b). Recently, research on language agents focus on agents’ ability to use language models for perceiving, reasoning, planning, and managing memory while interacting with external environments (Sumers et al., 2024). These agents, which learn independently from their observations and adapt their knowledge, are equipped with external storage for long-term memory, allowing them to store past reasoning (Majumder et al., 2024; Wang et al., 2023c; Zhao et al., 2024; Park et al., 2023; Shinn et al., 2023) or user feedback (Madaan et al., 2022) for future use and self-reflection (Madaan et al., 2024).

### v2

@@ 5.2. Consumption
In REML, the predictive model is presented with one or more documents. The effectiveness of $f_{\theta}$ is influenced by the consumption of the presented documents. Ideally, $f_{\theta}$ would consume all the documents simultaneously, yet our systems are computationally limited; hence, the granularity of consumption is typically limited to a subset of the presented documents. Depending on the task, different consumption algorithms may prove varying in utility—some algorithms are used for extraction and others for on-the-fly updates of the predictive model parameters. In contrast, decoding algorithms, such as beam search (Freitag and Al-Onaizan, 2017) or nucleus sampling (Holtzman et al., 2020), provide ways to decode effective outputs given the presented documents and can incorporate verification for improved effectiveness. There are additional concerns during consumption, such as efficiency (De Jong et al., 2023; Hofstätter et al., 2023) and attribution (Gao et al., 2023b; Schuster et al., 2024; Asai et al., 2024; Menick et al., 2022), that provide further utility. Table 3 summarizes the research related to consumption.
@@ 5.2.1. Consumption at different granularities of retrieval items
  • Joint: When the prediction has a sufficiently flexible context representation, then multiple retrieval items can be passed simultaneously in a single inference procedure (Izacard and Grave, 2021b; Lewis et al., 2020). This is potentially richer than the ensemble approach since each retrieved item is aware of the others. Due to computational limits, only a few retrieval items are able to be processed this way and the ensemble-approach is relatively more scalable.
@@
These paradigms are atomic functions that can be composed to form more complex operations. For instance, given lists of retrieved items, denoted as $r_{0},r_{1},r_{2},r_{3}$, the atomic functions with different granularities can be composed as:
@@ 5.2.2. Consumption algorithms
  • Extractive: The predictive model is limited to extracting exact information from the retrieved item, e.g., a span of text from a retrieve passage to answer a question (Khandelwal et al., 2020; Lan et al., 2023; Wang et al., 2023a). This can be achieved through pointer networks (Vinyals et al., 2015), constrained decoding (Hokamp and Liu, 2017; Hu et al., 2019; Post and Vilar, 2018), and other similar techniques (Kim et al., 2023).
@@ 5.2.3. Decoding algorithms
Independent of the consumption algorithm, there are different decoding algorithms that can be used for producing outputs in REML. The decoding algorithms explored thus far fall into the following categories:
@@
  • Retrieval-enhanced: Search algorithms like beam search will consider both the model output and the retrieved items when scoring candidate predictions (Lewis et al., 2020; Asai et al., 2024). This should penalize spurious associations between query and retrieved item.
@@ 5.2.5. Attribution and other extensions
Advanced applications of REML will augment the predictive model output space to incorporate REML-specific information. The most common instance of this is probably to support attribution, so that each part of the output can be traced back to the relevant retrieved item (Gao et al., 2023b; Schuster et al., 2024; Asai et al., 2024; Menick et al., 2022). Other cases involve in-line verification or calls to external tools that would not be easily possible without incorporating retrieved items (Asai et al., 2024).
Consumption Granularities (§5.2.1) $k$NN-LM (Khandelwal et al., 2020) Single + Ensemble Probabilities are computed for each retrieved item independently then combined. RePLUG (Shi et al., 2024) Single + Ensemble Probabilities are computed for each retrieved item independently then combined. RAG (Lewis et al., 2020) Joint Retrieved items are concatenated before consumption. FiD (Izacard and Grave, 2021b) Joint Retrieved items are concatenated in the decoder during consumption. FLARE (Jiang et al., 2023) Multi-round Potentially retrieves new items as generation progresses. Consumption Algorithms (§5.2.2) kNN-LM (Khandelwal et al., 2020) Extractive A single word is selected from the retrieved context. CoG (Lan et al., 2023) Extractive An extension of kNN-LM that can output both words and phrases. FilCo (Wang et al., 2023a) Extractive Filters out irrelevant texts from retrieved documents. CBR (Das et al., 2020) Analogical Incorporate knowledge graphs into neural models in spirt of case based reasoning. Dynamic L2M (Drozdov et al., 2023) Analogical Retrieves demonstrations for few-shot prompting. RePLUG (Shi et al., 2024) Contextual Uses retrieved items in the context. SPALM (Yogatama et al., 2021) Latent Incorporates retrieved items into the hidden state. Decoding Algorithms (§5.2.3) kNN-LM (Khandelwal et al., 2020) Text-only Retrieval probability is ignored for next word prediction. RAG (Lewis et al., 2020) Retrieval-enhanced Retrieval probability is incorporated in beam search. Self-RAG (Asai et al., 2024) Retrieval-enhanced Critic probability is incorporated in beam search. FLARE (Jiang et al., 2023) Verification-based Discards low confidence continuations, triggering retrieval. Consumption Efficiency (§5.2.4) LUMEN (De Jong et al., 2023) Precompute Partially computes passage representations offline. Attribution and Extensions (§5.2.5) ALCE (Gao et al., 2023b) Attribution End of sentence. Multi-source. SemQA (Schuster et al., 2024) Attribution Mid-sentence. Multi-source. Self-RAG (Asai et al., 2024) Attribution End of sentence. Single source. Self-RAG (Asai et al., 2024) Verification Outputs a special token indicating whether to use document.
@@ 6. Storing
Storing is one of the optional yet crucial properties of REML models and refers to how retrievable items are saved, represented, and indexed. The storage components can be categorized into coupled and decoupled storage. If at least one external memory is optimized jointly with the predictive model, we call the architecture has coupled storage. In the coupled storage architecture, contents can be populated to the storage online, updated alongside a predictive model, e.g., Neural Turing Machines (NTM) (Graves et al., 2014, 2016) and REALM (Guu et al., 2020). On the other hand, if all external storage are from off-the-shelf system, and the contents are populated offline, we call the architecture has decoupled storage, e.g., kNN-LM (Khandelwal et al., 2020) and ED2LM (Hui et al., 2022).
In its simplest form, REML systems will operate with decoupled storage where the retrieval model is implemented independent of the predictive model. Since entries in the storage are populated offline and many off-the-shelf retrieval models are readily available, it is relatively convenient to construct the decoupled storage. On the other hand, many advanced REML systems operate with coupled storage where the retrieval model is directly influenced by the predictive model. Here, the entries in the storage are populated online and updated alongside the predictive models. In this section, we describe storage operations and challenges associated with the coupled and decoupled storage.
@@ 6.1. Primary Storage Operations
In general, there are three types of operations that each storage system must support to be effectively utilized in REML systems: Address Generation, Read, and write. The storage operations will typically be used in three scenarios: 1) $f_{\theta}$ needs to incorporate historical context in its prediction, e.g., long context language models (Bertsch et al., 2023); 2) $f_{\theta}$ conducts various types of online learning using recent/past interactions, e.g., experience replay in reinforcement learning and language agents (Shinn et al., 2023); and 3) $f_{\theta}$ is a memory network-like architecture (Weston et al., 2015) where retrieval is an abstract process and everything written to or read from the storage is controlled by the network which is the service of the objective being optimized.
@@ 6.1.3. Write
where payload can be a form of vector or raw representation, which can be preprocessed by some functions, similar to the functions defined in Section 5, before being stored in the storage (Hui et al., 2022).
@@ 6.2.1. Storage Construction
REML systems mostly initialize its storage by processing an entire retrieval corpus. This can be done offline before training, after training before inference, and throughout training as needed. Storage construction is well documented and studied, and the initial storage construction is essentially a series of write and address generation operations. Storage can be constructed as a key-value structure where retrieval space $\mathcal{D}$ can be defined as:
@@
where $transform_{k}$ is a key representation function that can take each entry in the corpus or an input instance $x$. Similarly, $transform_{v}$ is a value representation function that can take each entry in the corpus or an input instance $x$ to generate a value in the storage. Note that the $transform_{k}$ and $transform_{v}$ function can simply be the identity function, meaning it does not change the key and value at all. Table 4 describes how each paper constructed its storage offline and/or online. For example, in EMAT (Wu et al., 2022d), where the collection is pairs of questions and answers, the $transform_{k}$ function utilizes only the question in each pair to generate the key representation, and $transform_{v}$ function uses only the answer from each pair to generate the value representation.
@@ 6.2.2. Storage Management
Once the storage is initialized, both for optimal task completion and efficiency, there is a need to schedule the storage operations, and we dub this phase as storage management. Efficient usage of storage can be understood in terms of space and speed which come down to: when, what, and how to store.
@@
In many scenarios, $g_{\omega}$ will pull information from the storage built at an initial phase based on the need of $f_{\theta}$. However, this setup can introduce storage staleness problem when $\omega$ changes during $g_{\omega}$ is being sequentially or jointly trained with $f_{\theta}$. Since $\mathcal{D}_{\text{key}}$ (retrieval space of keys in C) is constructed by $\omega$ during the training, it is necessary to refresh the storage as $\omega$ is updated; this was even mentioned from MemNN (Weston et al., 2015). This comes down to asking two questions: 1) how often to update, and 2) what portion of storage to update. For the first question, one can either synchronously (every training step) or asynchronously (every $T$ training step) update the storage (Asai et al., 2023a). For the second question, one can choose to update the entire storage, a subset of the storage, or refrain from updating the storage at all.
@@
This encompasses entry representation, e.g., index compression (Wu et al., 2022b; Rae et al., 2020; Martins et al., 2022) and quantization (Izacard et al., 2024) and architectural choice of the storage, e.g., key-value structure (Grave et al., 2017; Khandelwal et al., 2020; Wu et al., 2022c; Zhong et al., 2022; Min et al., 2023; Borgeaud et al., 2022; Yogatama et al., 2021; Alon et al., 2022; Hui et al., 2022), where the compression and representation computations can happen incrementally by batch (Zamani et al., 2022).
@@ 6.3. Storage Types
Coupled Storage (§6.3.1) Key Value NTM (Graves et al., 2014, 2016) transformed output of $f_{\theta}$ (the same as the Key) MemNN (Weston et al., 2015), MemN2N (Sukhbaatar et al., 2015) & DMN (Kumar et al., 2016) input feature embedding (the same as the Key) Neural Cache Model (Grave et al., 2017) hidden representation of RNN next word RUM (Chen et al., 2018) user-item embedding (the same as the Key) Transformer-XL (Dai et al., 2019) hidden representation of Transformer (the same as the Key) LongMem (Wang et al., 2023b) & Memorizing Transformer (Wu et al., 2022c) attention-key attention-value MemTransformer (Burtsev et al., 2021) sequence of tokens tokens RPT (Rubin and Berant, 2023) & Unlimiformer (Bertsch et al., 2023) token chunk embedding (the same as the Key) MeMViT (Wu et al., 2022b) & PFMN (Lee et al., 2018) image frame embedding (the same as the Key) REALM (Guu et al., 2020), REPLUG LSR (Shi et al., 2024) & ATLAS (Izacard et al., 2024) document embedding (the same as the Key) EMAT (Wu et al., 2022d) question embedding answer embedding QAMAT (Chen et al., 2023b) question embedding question-answer embedding TRIME (Zhong et al., 2022) context (sequence of tokens) next token NPM (Min et al., 2023) token embedding token Reflexion (Shinn et al., 2023), CLIN (Majumder et al., 2024) & ExpeL (Zhao et al., 2024) self-reflection in NL (the same as the Key) Generative Agent (Park et al., 2023) stream of experience in NL (the same as the Key) Voyager (Wang et al., 2023d) program description embedding program code MemPrompt (Madaan et al., 2022) NL question NL human feedback Synapse (Zheng et al., 2024) task metadata embedding state and trajectories in NL Decoupled Storage (§6.3.2) kNN-LM (Khandelwal et al., 2020) & SPALM (Yogatama et al., 2021) context embedding next token RAG (Lewis et al., 2020) document embedding (the same as the Key) RETOMATON (Alon et al., 2022) context embedding next token, pointer KIF (Fan et al., 2021) evidence embedding (multimodal) (the same as the Key) RETRO (Borgeaud et al., 2022) token chunk embedding token chunk ED2LM (Hui et al., 2022) document embedding document REPLUG (Shi et al., 2024) document embedding (the same as the Key) Embodied-RAG (Xie et al., 2024) node of a topological map in NL position and path to the location
@@
In the literature, two types of storage architectures are identified in REML systems: Coupled Storage and Decoupled Storage. The following sections and Table 4 will elaborate on these architectures.
@@ 6.3.1. Coupled Storage
Coupled storage is defined as a storage that can be updated online during training and inference of the predictive model and can be jointly optimized. Initial developments of the coupled storage enhancements were primarily led by Neural Turing Machine (NTM) (Graves et al., 2014, 2016) and Memory Network (MemNN) (Weston et al., 2015). They showed abstract operations, such as copying, recall, and sorting, to language reasoning tasks by leveraging external addressable storage, where contents in the storage of these architectures are dense vectors which can be readily used by $f_{\theta}$. We refer the readers to the original papers of the two models for deeper understanding of primary shape of REML models with coupled storage. There are a few notable characteristics about REML with coupled memory, including but not limited to the following.
@@
One of the primary concerns of the coupled storage arises when $g_{\omega}$ is being trained, making the storage stale. It is still an ongoing challenge in the community, but as mentioned in Section 6.2, there have been several techniques devised to circumvent this issue by answering how often to update (synchronous or asynchronous) and what portion of the storage to update (full or partial). Therefore, there are five different strategies including an avoidance of update. Synchronous full update is the simplistic approach to solve the staleness problem by updating the storage at every training step. It is attempted in a few research (Rubin and Berant, 2023; Bertsch et al., 2023), but its large computational overhead prevents it from being used in a practical setting (Izacard et al., 2024). Synchronous partial update can be done by selecting a batch of entries to update (Izacard et al., 2024). Depending on the applications, there can be various batch selection strategies, such as lexical similarity (Zhong et al., 2022) or in-document sampling (Min et al., 2023). Asynchronous full update is done by updating the full storage every $T$ training steps (Guu et al., 2020; Izacard et al., 2024; Shi et al., 2024; Wu et al., 2022d). This allows staleness in the index before it is updated again. For example, Wu et al. (2022d) freeze the storage at the beginning of each training epoch and only updates at the end of each epoch. As far as we know, there is little attempt on asynchronous partial update as it may degrade the training performance in larger margin. Alternatively, since index recreation is highly expensive, it is possible to ignore the staleness and avoid re-indexing with adequate strategies without compromising a performance in a large margin (Rae et al., 2016; Lewis et al., 2020; Izacard et al., 2024; Guu et al., 2020; Wang et al., 2023b; Wu et al., 2022c).
@@
Another characteristic of the models equipped with coupled storage is that they can have a cold start problem, where the performance of $f_{\theta}$ is suboptimal before the storage is filled up with enough information. Most of the architectures that start with an empty storage such as language agents (LA) or long-context language models that process a long document in multiple training steps have this issue (Park et al., 2023; Majumder et al., 2024; Shinn et al., 2023; Zhao et al., 2024; Wang et al., 2023d; Wu et al., 2022c; Zhong et al., 2022). However, the cold start problem can be alleviated when the model can be adapted to a new task and the storage/experience built during the previous tasks is transferable to the new task (Majumder et al., 2024).
@@
Despite its disadvantages, coupled storage has seen significant development both theoretically and in various applications. It includes making REML systems with coupled storage end-to-end trainable (Sukhbaatar et al., 2015), capturing the position and temporality of language using episodic storage (Kumar et al., 2016), and scaling the storage through sparse access (Rae et al., 2016; Zaremba and Sutskever, 2016) and strategic storage management (Lample et al., 2019; Grave et al., 2017). These have led to applications in various domains, such as meta-learning (Santoro et al., 2016), sequential recommendation (Chen et al., 2018), video summarization and recognition (Lee et al., 2018; Wu et al., 2022b). In long context language modeling (Bertsch et al., 2023; Zhong et al., 2022), there have been approaches to solve the task through attention recurrence (Dai et al., 2019; Yogatama et al., 2021; Wu et al., 2022c; Wang et al., 2023b) and compression of hidden states (Rae et al., 2020; Martins et al., 2022; Wu et al., 2022b). Recently, research on language agents focus on agents’ ability to use language models for perceiving, reasoning, planning, and managing memory while interacting with external environments (Sumers et al., 2024). These agents, which learn independently from their observations and adapt their knowledge, are equipped with external storage for long-term memory, allowing them to store past reasoning (Majumder et al., 2024; Wang et al., 2023d; Zhao et al., 2024; Park et al., 2023; Shinn et al., 2023) or user feedback (Madaan et al., 2022) for future use and self-reflection (Madaan et al., 2024).